OpenStudy (anonymous):

There are 23 coins in a bank. All the coins are dimes and quarters. The total value of the coins is $3.80. How many dimes are there? How many quarters?

OpenStudy (australopithecus):

You have two formulas where dimes are y and quarters are x we have the formula 0.25(x) + 0.10y = $3.80 and another formula x+y = 23

OpenStudy (australopithecus):

substitute one equation into the other to solve for one of the variables then sub that variable back into one of the original formulas to solve

OpenStudy (anonymous):

Set up a system of equations. \(\ \huge x+y=23 \). \(\ \huge .25x + .10y = 3.8 \). Now solve for x and y.

OpenStudy (australopithecus):

so set x = 23 - y 0.25(23-y) + 0.10y = $3.80 now solve for y once you have solved for y sub it back into one of the original equations to solve for x

OpenStudy (australopithecus):

doing this tells us where the graphs intersect and therefore the answer we are interested in